Tech Staff Meeting Notes September 9, 2016

The top priority is always repair of down stations.
NM problem stations
ET problem stations

Duty Table updated June 2014


News & Highlights

Jim and Dave will be attending NetOpsVIII in Pasadena on October 11-12. More info here:

SIS training session is set for Thursday 9/15.

The CERI General Meeting (aka Chuck's State of the Union Address) is scheduled for Monday 9/19 3-4:30.

Primatol is condisered a Hazardous Chemical so we need ES&H to approve an exemption at purchasing.


We have an obsidian to replace the CMG5TD at NMEM.


Data taken from: /gaia/smeagol/dat2/Working/NM/201609/20160903_120300_MAN

NM Down

NM Misc

AG Misc
LCAR.HN*- should see a signal on this station for the 5.8 Oklahoma quake but there is nothing.

PC at NHIN will not power up. Steven will get a replacement ready. Can install the obsidian at NMEM on the way.

WYBT was vandalized for the second year in a row. Chris got it going temporarily. Need to revisit and harden.

MORT is closed. Need to look for a compromise site to fill the hole left by it and HAWT. Needs to assume flooding. Not a top priority at this time.

Have three of the five remaining ISIS3 upgrades for the NMAD node are complete. We have hardware on deck fo the last two and another three ready to go to the MKTA node.

Sensor upgrade on the I40 Bridge is is complete. Cable for D2 sensor is bad. Vertical on the G7 borehole is also odd with a large vertical offset. There may be a few other channels that need a second look. A visit is scheduled for Wednesday 9/14.

Also, it turns out that we've been archiving 200 sps data as HDBR2 since switching to the Granites but we have to use a single station name for both 100 and 200 sample rate datastreams. Currently the metadata only use the 100 sps channels and have an ondate in September 2012 (e.g. pre-granite epoch doesn't exist). So all existing HDBR2 data needs to be converted to HDBR. All the 100 sps data will use location codes where the first character numberically represents the DAS it came from (i.e. 1, 2, 3) and the second character the sensor. The 200 sps data will use a letter for the DAS and a number for the Sensor. So channel A1 is 200 sps and 11 is 100, and the channel with the B1 location code is 200 sps and the same sensor a 100 has location code 21. Mitch will fix the metadata and Steve will correct the archived data.

There was an email discussion on the history of OXF which led to digging up the MOU from 1993. Mitch checked with Dave Wilson at ASL and it may be possible to replace the MOU, which is endorsed at a fairly high level within USGS, with a "station agreement" which is less formal.

NODE Tasks

Steve asked about replacing all node disks with solid state disk. Funds are available. Steve will get an estimate.

All nodes are due a firmware upgrade that will sufficiently change things such that the the initial upgrade and configure should be done at CERI, then swap hardware at the node.

John checked other networks plans to replace Windows digiziting boxes and found none. NC has migrated to the NI X-series boards and adsend does work with it though its still on MS Windows. John will try to find out which specific board and OS is used.

DBB: Move the DBB link at CERI from the west end of 3892 to the east side of 3904 and eliminate the thin-wire. Tower is up to test the waters.



Uverse is likely available at the lnxt node. Steve will investigate to see if its cost effective to upgrade from dsl. We don't need the band width but service for dsl will likely continue to deteriorate.




Flooding in the expansion space has become a problem. A Workorder was submitted and P-plant engineers were on-site to assess the problem. The most likely solution will be a retention wall and removal of material to make the foundation above grave.


We discovered that most of the ISIS4 packages have a wire reversal that inverts the orientations. All but two ISIS4 stations are impacted (RCGA and DSTN). John will visit and repair each station. John will also create a spreadsheet that lists each station, when the ISIS4 was intalled and when it was repaired so that Mitch can correct the orientations in the metadata.

Replacement station for GTTN and RBNC being evaluated and permissioned for Southern Highlands Science Learning Center at Purchase Knob, a part of the Great Smoky Mountains National Park. Noise study picked up compressor noise from a nearby air quality station. Need to add isolation mounts and test farther away. We had to return the compact so will need to use a 6TD instead. Cell service is probably not available so a non-real-time 6TD is needed.


Taken from: /gaia/smeagol/dat2/Working/ET/201609/20160907_000700_MAN

Misc Problem

GRBT.EHE see image
FPAL.HHE - drifting? see image

Drifting at FPAL is getting worse. A CERI crew will service PWLA then continue on and replace the 120PA at FPAL with a 120AQ. Mitch with check with Nanometrics if its cost effective to upgrade the PA to a QA.

GMG has had a power upgrade, new batteries, new solar contoller, and better surge suppression. And it still went down on 1/18. James can get to the radio at the station which is on a separate power feed than the reftek. Jim will send a new BC10 and John will investigate.

CVVA needs to be DHCP instead of static IP. James will configure the fanless pc as a miniworm which will run an export_actv to iniate the connection from there to here. This means we still need an static ip, but it can be a 192 behind a router and the router can be DHCP to the outside world.

Need to return the bad 5TD from SEAR for an obsidian (to be installed at CVVA).

John's updated todo list .


Improve RCGA AC power and perform maintenance work on shed.

Move PMTN due to excessive noise and vandalism. The trees at Lone Mountian are too tall to get a signal out. Looking for an alternate site.


HPNC node is up. Need to fix Win7 to allow earthworm commands (e.g. status, restart etc) from the ssh command line. 9 channels(GFM, MGNC, and DSNC) are being exported over a cell modem temporarily until internet is established. Still waiting on completion of communcations to get on local internet. Its still on a cell modem; Randy Loop plans to install the internet modem next week. Hopefully after that we can retrieve the cell modem and get all channels to CERI in real-time.

Upgrade remaining High Peak Node Stations from L4-3D's to S-13's. At least one bad component in L4-3's at CCNC, TVNC, TRYN.

GFM2 is good to go. It will be a reftek with S-13's. Do we need to take student volunteers to help with the GFM decommission?


WMTN pit needs to be moved due to erosion.

CPRT comm reconfig. John says, "CPRT is going through the Senior Center at Corryton (cort). We removed everything from svfd an everything but what is on the tower at gttn. cort is indeed using dynamic addressing but data is flowing. Greg Hill the director at the senior center is supposed to be contacting his IT person to see what can be done. I left a message for Greg last week but still haven't heard back. I want to get Greg's IT guy in direct contact with either Steve or James so they can figure it all out."


Bob has funding for a new CUS station. Deadline is Aug 31 for obligation of all funds. It will be a drill braced monument rather than the piling style. MODOT already has a good quality station near Hayti so Chris is now looking at Ridgely. Site A is a big NO. Need to continue the search.

CVMS has intermittent connection problems with wireless. Its hardwired for now but the wireless needs to be replaced to preserve electical isolation between the building and vault.

All GPS stations are back up.

AG Network

Z41A is online with AG. Vault leaks so did not install sm sensor. Need to fix the leak and install the sm vault inside the TA vault.

Pulled trillium/reftek from X702. Need to install an AGS 6TD in its place. Done.

When the water leak at LCAR was repaired, was an accelerometer reinstalled?


Needs some clean up.

The FDL wiki is here:


The seisnet group on umdrive is at You will need your uuid and password.

The seisnet wiki is at You will need your uuid and password.